OpenStudy (anonymous):

1OVER 4X -1 OVER 3 =1 OVER 2

OpenStudy (saifoo.khan):

\[\frac{1}{4x}- \frac{1}{3}=\frac{1}{2} \] this?

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

he's just trolling

OpenStudy (saifoo.khan):

Who He?

OpenStudy (anonymous):

ralph

OpenStudy (saifoo.khan):

Damn it.

OpenStudy (anonymous):

lol its ok, got pranked by hashir today

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

Hi Book.

OpenStudy (saifoo.khan):

x = 3/10

OpenStudy (anonymous):

saifoo u r bhad

OpenStudy (saifoo.khan):

bhad? o.O

OpenStudy (callisto):

1/4x -1/3 =1/2 (3/12x) - (4x-12x) = 1/2 6-8x = 12x 20x=6 x=3/10

OpenStudy (anonymous):

1/4x -1/3 =1/2 (3/12x) - (4x-12x) = 1/2 6-8x = 12x 20x=6 x=3/10

OpenStudy (anonymous):

oh callisto i think i plagiarised